Université de Lyon - Université Lyon 1

Examen terminal du 12 janvier 2009

Espaces de Hardy et applications/ Groupes algébriques linéaires

Aubry Cédric Bayard Médéric Chabert Loïc Chatard Marianne Ducet Virgile Ducort Jérôme Guillaud Nicolas Hoang Duc Auguste Iosti Simon Lassoued Dahou Lebre Marie-Ange Liboz Emilie Maaref Walid Male Camille Motier Camille Rey Thomas

2